x2 + x +6 The value of the constant k such that lim exits is X-1 Select one: -5 - 12 5 -6 -7
Answer #1

22.2250 7/8 5/8 = 15,8750 9.5250 3/8 = 3.1750 1/8 = POPU Page No. 3/4 = 19.0500 6.3500 1/4 = The value constant k K such 12.7